Koontz Company manufactures a number of products. The standards relating to one of these products are shown below, along with actual cost data for May Standard Cost per Unit Actual Cost per Unit Direct materials $ 5.40 Standard: 1.80 feet at $3.00 per foot Actual: 1.80 feet at $3.30 per foot $ 5.94 Direct labor 16.20 Standard: 0.90 hours at $18.00 per hour Actual: 0.92 hours at $17.50 per hour 16.10 Variable overhead 4.50 Standard: 0.90 hours at $5.00 per hour Actual: 0.92 hours at $4.50 per hour 4.14 $ 26.18 $ 26.10 Total cost per unit Excess of actual cost over standard cost per unit $0.08 The production superintendent was pleased when he saw this report and commented: This $0.08 excess cost is well within the 2 percent limit management has set for acceptable variances. Is obvious that theres not much to worry about with this product. Actual production for the mh was 12 000 units Variable overhead cost is assigned to products on the basis of direct labor-hours There were no beginning or ending inventories of materials Required: 1. Compute the following variances for May a Materials price and quantity variances. b. Labor rate and efficiency variances. c Variable overhead rate and efficiency variances. 2. How much of the $0.08 excess unit cost is traceable to each of the variances computed in (1) above 3 How much of the $0.08 excess unit cost is traceable to apparent inefficient use of labor time?
1a. Compute the following variances for May, materials price and quantity variances. 1b. Compute the following variances for May, labor rate and efficiency variances. 1c. Compute the following variances for May, variable overhead rate and efficiency variances. (Indicate the effect of each variance by selecting F for favorable, U for unfavorable, and None for no effect (i.e., zero variance). Input all amounts as positive values.) Show less 1a. Materials price variance Materials quantity variance 1b. Labor rate variance Labor efficiency variance 1c. Variable overhead rate variance Variable overhead efficiency variance Required 1 Required 2
Answer #1

Solution 1a:

Material price variance = (SP - AP) * AQ = ($3 - $3.30) * 12000*1.80 = $6,480 U

Material quantity variance = (SQ - AQ) * SP = (12000*1.80 - 12000*1.80) * $3 = 0

Solution 1b:

Labor rate variance = (SR - AR) * AH = ($18 - $17.50) * 12000*0.92 = $5,520 F

Labor efficiency variance = (SH - AH) * SR = (12000*0.90 - 12000*0.92) * $18 = $4,320 U

Solution 1c:

Variable overhead rate variance = (SR - AR) * AH = ($5 - $5.50) * 12000*0.92 = $5,520 F

Variable overhead efficiency variance = (SH - AH) * SR = (12000*0.90 - 12000*0.92) * $5 = $1,200 U

Solution 2:

Statement showing Excess unit cost traceable to variance - Koontz Company
Variance
Materials:
Price Variance $6,480.00 U
Quantity Variance $0.00 F $6,480.00 U
Labor:
Rate variance $5,520.00 F
Efficiency variance $4,320.00 U $1,200.00 F
Variable overhead:
Rate variance $5,520.00 F
Efficiency variance $1,200.00 U $4,320.00 F
Excess of actual cost over standard cost per unit $960.00 U

Solution 3:

Statement showing Excess unit cost traceable to variance
Variance
Excess of actual over standard cost per unit $960.00 U
Less: Portion attributable to labor efficiency:
Labor Efficiency variance $4,320.00 U
Variable overhead Efficiency variance $1,200.00 U $5,520.00 U
Portion due to other variances $4,560.00 F